# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

### v2

@@ 3 Augmenting the Action Space with Conventions
Similar to how options enabled the incorporation of existing domain knowledge into RL algorithms by modifying an agent’s action space, we propose a MARL solution strategy to achieve a similar goal. Options required the defining of semi-MDPs, where actions can have varying lengths, and in a MARL scenario will theoretically require decentralised partially observable semi-MDPs (encapsulated by Amato et al.’s MacDec-POMDP [46]). However, our solution strategy does not directly change the length of an action by having an external policy take over, rather it facilitates implicit communication through a specific behaviour as a result of a sequence of actions, based on predefined principles, called a convention. Each convention contains convention-steps that span over and include multiple time steps and multiple agents, requiring them to actively opt in for the convention to reach fruition. Therefore, our augmentation of the action space does not directly change the Dec-POMDP framework, and merely acts as an extension of the existing action space to allow for more advanced behaviours.
Ideally an agent’s action space can be comprised of a pure-convention space, i.e. containing only convention-steps and no primitive actions. However, there can exist scenarios where no convention-steps are available, and the agents will not be able to take any actions. To solve this problem, we propose augmenting the action space with primitive actions and conventions simultaneously, similar to the discussion by Sutton et al. [19] on combining primitive actions and options at the slight cost of performance. We will prove in Section 4.3 at the hands of the Small Hanabi problem, that this solution strategy only slightly impacts the overall performance when comparing agents with a pure-convention space and an augmented action-convention space.
@@ 4 Hanabi
To test the capabilities of conventions, we will use the cooperative Hanabi environment, which was proposed as a frontier for MARL algorithm development by Bard et al. [36]. Hanabi is a 2–5 player card game, best described as a cooperative solitaire [36]. Players have a hand of five cards for two and three player, and four cards for four and five player. Each card has a suit/colour (red, yellow, green, white, or blue) and a rank (1 to 5). The deck comprises 50 cards total, with 10 cards for each suit with a distribution of three 1’s, two 2’s, 3’s, and 4’s, and finally only one 5. The aim of the game is to stack these suits in numerical order from 1 to 5, however players cannot see their own cards (i.e. their hands are hidden), but can see the cards of their fellow players.
@@
Figure 2: An example game of Hanabi as seen from the perspective of player 1. There are four hint tokens left and the players have lost one of their shared life tokens. Player 1 knows about two 4s in their hand and the green, blue and yellow stacks have been partially completed, leading to a current game score of 6/25. It is now player 1’s turn, and they can take the hint (to player 2 or 3), play (from their hand) or discard (from their hand) action.
@@ 4.2 Artificial Conventions in Hanabi
Conventions Example 2. In Fig. 2 each player is controlled by an agent, with the current player (referred to as the active player) being agent 1. The convention at hand (as mentioned in Conventions Example 1), which we will refer to as $c_{1}$, states that if an ambiguous hint is given, the focus of that hint is the left-most (or newest) card. Since the condition $\lambda_{1}^{1}$ is met, i.e. player 2 has two cards in their hand that share the same colour or rank and the left-most one is playable, the action {start $c_{1}$} is available for agent 1. Agent 1 chooses to initiate $c_{1}$ which in turn gets translated by the policy $\pi_{1}^{1}$ to an environment action of {hint red to player 2}, ending player 1’s turn. It is now player 2’s turn, and they have just received an ambiguous colour hint from player 1, and given their current observation they can derive that $c_{1}$ is currently active. This triggers the subscribing (and also completing) condition $\lambda_{1}^{2}$ (it is completing since $m_{1}=2$), where agent 2 can now opt in to continue (and complete) $c_{1}$. When agent 2 chooses to subscribe to $c_{1}$, their action is translated by the policy $\pi_{1}^{2}$ to an environment action of {play card in position 1}, subsequently ending player 2’s turn and completing $c_{1}$.
@@ 4.3 Preliminary Results on Small Hanabi
To test the viability and capabilities of conventions, we conduct initial tests using the Small Hanabi environment, since it offers a smaller state-action space and allows for a pure convention space [^12] A pure convention space contains only conventions, i.e. no primitive actions, and accounts for all possible scenarios, ensuring the agents will always have a convention available. In Small Hanabi this is possible due to the small state-action space, but becomes a far greater challenge in the full Hanabi problem.. We compare a pure convention space and an augmented action-convention space, as discussed in Section 3.2, against their primitive action benchmarks. Due to Hanabi having a long list of intricate conventions, in addition to augmenting the action space with conventions, we experiment with the idea of simplifying the list of conventions during augmentation [^13] This idea stemmed from the discussion by Sutton et al. [19], where they explored the capabilities of combining primitive actions with simplified options, since the primitive actions can account for the simplification at the slight cost of performance.. This simplified list consists of the conventions that were the most straightforward and natural to develop, and were also the most common ones to occur in a game. The removed conventions can be considered edge-case or specific to certain game states that are very uncommon, and often results in uncertainty regarding their effectiveness.
Small Hanabi has a deck reduced to 20 cards with only two suits, is limited to only two players, and the player hand size has been reduced to two cards. Furthermore, the life tokens have been reduced to one, and the hint tokens to three, otherwise the game shares all the core mechanics as regular Hanabi. This results in the problem having a significantly smaller state-action space and allows for initial testing and development of algorithms, with the aim of solving the full Hanabi problem. Even though this problem is considered simpler than Hanabi, it is by no means an easy problem to solve, with a perfect score of 10/10 being difficult to achieve. Player hands are often “locked” with important cards that must not be discarded (called critical cards), hint tokens run out almost immediately, and players are often forced into no-win scenarios where they must choose a bad action or risk losing their single life token.
@@ 5 Performance Evaluation Using Hanabi
The preliminary tests for action augmentation with conventions in Small Hanabi has shown promising results, and allow the agents to achieve faster training time and improved policies. We now shift focus to the full Hanabi problem for 2–5 players and with the conventions defined by the H-Group and discussed in Appendix A.2. Our results will focus on Bard et al.’s [36] Rainbow agent as a baseline, and apply an augmented action-convention space to improve on its performance. Rainbow was chosen as our baseline since it is among the methods that train the fastest, is the most sample efficient among the existing Hanabi agents [36, 13, 14, 40, 27], has little run to run variance [36], is considered a widely applicable algorithm [57, 58, 59], and still leaves considerable performance to be desired within Hanabi. All tests are conducted on the open-sourced Hanabi learning environment developed by Bard et al. [36].
@@ 5.1 Experimental Setup
Before discussing the results, we will briefly highlight the architecture design for each deep RL technique. For a full list of the hyperparameters used in each method see Table 6 in Appendix B. All methods receive a one-hot encoded observation defined by the environment as input to their neural networks, notably (by default) this observation tuple contains the most recent actions within the previous round, and is not added as an additional feature. The Rainbow agents use a Multilayer Perceptron (MLP) consisting of two feed forward neural networks with 512 neurons each, with the ReLU activation function [60] applied. It implements the Adam optimizer [55] to calculate the TD error and perform backpropagations to update the network weights. Additionally, the Rainbow agents use distributional reinforcement learning to predict the value distributions which are approximated as a discrete distribution over 51 uniform atoms, along with prioritise replay memory sampling [28]. Even though Rainbow uses n-step bootstrapping, in these experiments a value of n=1 was found to be optimal, additionally noisy nets have been disabled in favour for a traditional decaying epsilon-greedy approach.
@@ 5.2 Self-play Performance
The learning curves for Rainbow with and without an augmented action-convention space for 2–5 player Hanabi are shown in Fig. 4. The two player scenario seen in Fig. 4(a) shows the smallest improvement when applying conventions, however the two player Hanabi problem can be considered a special case. This is due to the fact that the shared information between the two players are far more limited when compared to higher player counts. For example, in two player Hanabi, half of the hands are hidden to a single player whereas in the three player scenario only a third of the hands are hidden, and the players always have a common hand visible. This allows for more advanced reasoning over another player’s actions, evident by the fact that two of the conventions, specifically the Finesse and the Prompt as discussed in Appendix A.2, cannot be applied in two player Hanabi.
(a) 2 Player learning curves
(b) 3 Player learning curves
(c) 4 Player learning curves
(d) 5 Player learning curves
Figure 4: Learning curves with exponential moving averages (weight=0.9995) for Rainbow, obtained from Bard et al. [36], as baseline compared to Rainbow with an augmented action-convention space for Hanabi two to five players. The best agent is highlighted in black for each agent scenario within each player count.
Despite the fact that the complexity of the problem increases exponentially as more agents are added to the environment, the performance uplift of conventions become more apparent. In the three player setting seen in Fig. 4(b), the agents are able to achieve a significantly faster training time as well as an improved convergent performance with a higher average score. The performance improvement becomes even more apparent when looking at the five player setting, seen in Fig. 4(d), which is generally considered the most difficult problem to solve, and where the agents train roughly five times faster than the baseline Rainbow agents. In Fig. 5 the distribution of scores during evaluation of each agent over the course of 1000 episodes is shown. In each scenario, conventions allow for a significant performance improvement, achieving a consistent grouping near a score of 21/25 and less overall variance when compared to baseline Rainbow. Rainbow with an augmented action-convention space also displays a robustness to increased player counts, with a remarkably consistent behaviour across all scenarios.
(a) 2 Player
(b) 3 Player
(c) 4 Player
(d) 5 Player
@@
Table 1 shows the average score for the best agent from each player count’s various runs compared to the results from Bard et al. [36]. Rainbow with an augmented action-convention space demonstrates an improved score over baseline Rainbow for each scenario, with a higher average in the 3–5 player scenario and overall less deviation. Additionally, the three and five player results are competitive compared to other algorithms, such as ACHA, while also requiring substantially less training steps. In the case of five-player Hanabi, Rainbow with conventions outperform ACHA where ACHA required 20 billion training steps and Rainbow with conventions only required 30 million. A statistical significance test, conducted in Appendix C, shows that majority of the results are statistically different. However, the Rainbow with conventions agent’s scores are statistically similar to that of baseline Rainbow two players and ACHA three players, however conventions allow the Rainbow algorithm to reach these scores significantly faster (5x compared to baseline Rainbow and 1000x compared to ACHA). Furthermore, even though these two instances are statistically similar, the agents learn completely different policies, as seen by the cross-play performance evaluation in the next section.
@@ 5.3 Cross-play Performance
Table 2 shows the two player performance for various agents in cross-play Hanabi taken from existing literature, compared to the results of our cross-play Rainbow agents with an augmented action-convention space. The results for the Rainbow with conventions agents are statistically different from the other results reported in Table 2, as seen by Table 7 in Appendix C. Conventions are able to significantly improve on the performance of cross-play agents, allowing the baseline Rainbow agents to go from abysmal performance to an acceptable score. Furthermore, it is able to outperform SAD with only other-play applied for the two player scenario, with other-play and auxiliary task applied to SAD still holding the best score. However, as mentioned previously, auxiliary tasks only benefit the two player scenario in Hanabi [13]. When moving from self-play to cross-play, Rainbow with an augmented action-convention space only suffered a slight performance decrease and maintained remarkable consistency across various agent parings.
Table 2: Performance evaluation in two player Hanabi for the best performing agents in a cross-play scenario over the course of 1000 episodes. The Rainbow with an augmented action-convention space agents were chosen from the top performing two agents in 10 different runs. Each score is averaged out of 25 with the standard error of the mean shown in brackets.
@@
Table 3 shows the results for cross-play Rainbow with an augmented action-convention space for all player counts of Hanabi. As seen in the self-play results for Rainbow with conventions and confirmed in Table 3, the two player scenario is a special case, with the agents showing the biggest performance degradation when moving from self-play to cross-play, and overall worst score compared to the other cross-play results. While most research efforts on cross-play Hanabi focus on the two player scenario, our results demonstrate that more powerful conventions are applicable in higher player counts, thus, it is important to look at all the player scenarios when evaluating cross-play performance.
@@
Most remarkable is the five player results, where the cross-play Rainbow agents with an augmented action-convention space are able to achieve a relatively high score and still cooperate effectively. The reason conventions are able to offer such a large performance gain for cross-play is due to the fact that the agents learn from a communal list of conventions that are taken from existing domain knowledge. Thus, if an agent is paired with a never-before-seen cooperator that has learned to play with the same list of conventions, they will still be able to cooperate effectively, since there is no uncertainty over another agent’s intent and reasoning. This is not unexpected, since humans who learn from the same “list” of conventions (whether it is physical or passed down), and that have never before interacted, can cooperate and communicate effectively with relatively low uncertainty.
@@ 6 Conclusion
Our main results show that conventions are able to significantly improve on the training time for Rainbow agents in Hanabi, and in the case of three to five players, also improve on the converged policy. Other Hanabi algorithms, such as SAD or MAPPO, require billions of training steps to reach convergent policies, whereas conventions reduced the number of training steps for Rainbow to below 30 million consistently. Additionally, conventions showed significant performance increase for cross-play agents, and allowed the Rainbow agents to go from not cooperating at all to being able to achieve decent performance across all player scenarios. For the most difficult scenario of five players, the cross-play Rainbow conventions agents trained on 50 million steps are able to outperform the ACHA self-play agents trained on 20 billion steps. Our work, as presented in this paper, mainly focused on single-and two-step conventions in the Hanabi problem, with the goal of extending these concepts to other multi-agent scenarios (including simultaneous action spaces) containing multistep conventions, as well as the exploration of more advance conventions (and the application of those conventions) on more advanced algorithms within the Hanabi setting.

### v3

@@ 3 Augmenting the Action Space with Conventions
Similar to how options enabled the incorporation of existing domain knowledge into RL algorithms by modifying an agent’s action space, we propose a MARL solution strategy to achieve a similar goal. Options required the defining of semi-MDPs, where actions can have varying lengths, and in a MARL scenario will theoretically require decentralised partially observable semi-MDPs (encapsulated by Amato et al.’s MacDec-POMDP [47]). However, our solution strategy does not directly change the length of an action by having an external policy take over, rather it facilitates implicit communication through a specific behaviour as a result of a sequence of actions, based on predefined principles, called a convention. Each convention contains convention-steps that span over and include multiple time steps and multiple agents, requiring each agent to actively opt in for the convention to reach fruition. Therefore, our augmentation of the action space does not directly change the Dec-POMDP framework, and merely acts as an extension of the existing action space to allow for more advanced behaviours.
Ideally an agent’s action space can be comprised of a pure convention space, i.e. containing only convention-steps and no primitive actions. However, there can exist scenarios where no convention-steps are available, and the agents will not be able to take any actions. To solve this problem, we propose augmenting the action space with primitive actions and conventions simultaneously, similar to the discussion by Sutton et al. [20] on combining primitive actions and options at the slight cost of performance. We will prove in Section 4.3 at the hands of the Small Hanabi problem, that this solution strategy only slightly impacts the overall performance when comparing agents with a pure convention space and an augmented action-convention space.
@@ 4 Hanabi
To test the capabilities of conventions, we will use the cooperative Hanabi environment, which was proposed as a frontier for MARL algorithm development by Bard et al. [37]. Hanabi is a 2–5 player card game, best described as a cooperative solitaire [37]. Players have a hand of five cards for two-and three-player, and four cards for four-and five-player. Each card has a suit/colour (red, yellow, green, white, or blue) and a rank (1 to 5). The deck comprises 50 cards total, with 10 cards for each suit with a distribution of three 1’s, two 2’s, 3’s, and 4’s, and finally only one 5. The aim of the game is to stack these suits in numerical order from 1 to 5, however players cannot see their own cards (i.e. their hands are hidden), but can see the cards of their fellow players.
@@
Figure 2: An example game of Hanabi as seen from the perspective of player 1. There are four hint tokens left, and the players have lost one of their shared life tokens. Player 1 knows about two 4s in their hand and the green, blue and yellow stacks have been partially completed, leading to a current game score of 6/25. It is now player 1’s turn, and they can take the hint (to player 2 or 3), play (from their hand) or discard (from their hand) action.
@@ 4.2 Artificial Conventions in Hanabi
Conventions Example 2. In Fig. 2 each player is controlled by an agent, with the current player (referred to as the active player) being agent 1. The convention at hand (as mentioned in Conventions Example 1), which we will refer to as $c_{1}$, states that if an ambiguous hint is given, the focus of that hint is the left-most (or newest) card. Since the condition $\lambda_{1}^{1}$ is met, i.e. player 2 has two cards in their hand that share the same colour or rank and the left-most one is playable, the action {start $c_{1}$} is available for agent 1. Agent 1 chooses to initiate $c_{1}$ which in turn gets translated by the policy $\pi_{1}^{1}$ to an environment action of {hint red to player 2}, ending player 1’s turn. It is now player 2’s turn, and they have just received an ambiguous colour hint from player 1, and given their current observation they can derive that $c_{1}$ is currently active. This triggers the subscribing (and also completing) condition $\lambda_{1}^{2}$ (the condition is completing since $m_{1}=2$), where agent 2 can now opt in to continue (and complete) $c_{1}$. When agent 2 chooses to subscribe to $c_{1}$, their action is translated by the policy $\pi_{1}^{2}$ to an environment action of {play card in position 1}, subsequently ending player 2’s turn and completing $c_{1}$.
@@ 4.3 Preliminary Results on Small Hanabi
To test the viability and capabilities of conventions, we conduct initial tests using the Small Hanabi environment, since it offers a smaller state-action space and allows for a pure convention space [^12] A pure convention space contains only conventions, i.e. no primitive actions, and accounts for all possible scenarios, ensuring the agents will always have a convention available. In Small Hanabi this is possible due to the small state-action space, but becomes a far greater challenge in the full Hanabi problem.. We compare a pure convention space and an augmented action-convention space, as discussed in Section 3.2, against their primitive action benchmarks. Due to Hanabi having a long list of intricate conventions, in addition to augmenting the action space with conventions, we experiment with the idea of simplifying the list of conventions during augmentation [^13] This idea stemmed from the discussion by Sutton et al. [20], where they explored the capabilities of combining primitive actions with simplified options, since the primitive actions can account for the simplification at the slight cost of performance.. This simplified list consists of the conventions that were the most straightforward and natural to develop, and were also the most common ones to occur in a game. The removed conventions can be considered edge-case or specific to certain game states that are very uncommon, and often result in uncertainty regarding their effectiveness.
Small Hanabi has a deck reduced to 20 cards with only two suits, is limited to only two players, and the player hand size is reduced to two cards. Furthermore, the life tokens have been reduced to one, and the hint tokens to three, otherwise the game shares all the core mechanics as regular Hanabi. This results in the problem having a significantly smaller state-action space and allows for initial testing and development of algorithms, with the aim of solving the full Hanabi problem. Even though this problem is considered simpler than Hanabi, it is by no means an easy problem to solve, with a perfect score of 10/10 being difficult to achieve. Player hands are often “locked” with important cards that must not be discarded (called critical cards), hint tokens run out almost immediately, and players are often forced into no-win scenarios where they must choose a bad action or risk losing their single life token.
@@ 5 Performance Evaluation Using Hanabi
The preliminary tests for action augmentation with conventions in Small Hanabi has shown promising results, and lead to the agents achieving faster training times and improved policies. We now shift focus to the full Hanabi problem for 2–5 players and with the conventions defined by the H-Group and discussed in Appendix A.2. Our results will focus on Bard et al.’s [37] Rainbow agent as a baseline, and apply an augmented action-convention space to improve on its performance. Rainbow was chosen as our baseline since it is among the methods that train the fastest, is the most sample efficient among the existing Hanabi agents [37, 13, 14, 41, 28], has little run to run variance [37], is considered a widely applicable algorithm [58, 59, 60], and still leaves considerable performance to be desired within Hanabi. All tests are conducted on the open-sourced Hanabi learning environment developed by Bard et al. [37].
@@ 5.1 Experimental Setup
Before discussing the results, we will briefly highlight the architecture design for each deep RL technique. For a full list of the hyperparameters used in each method see Table 6 in Appendix B. All methods receive a one-hot encoded observation defined by the environment as input to their neural networks, notably (by default) this observation tuple contains the most recent actions within the previous round, and is not added as an additional feature. The Rainbow agents use a Multilayer Perceptron (MLP) consisting of two feed forward neural networks with 512 neurons each, with the ReLU activation function [61] applied. It implements the Adam optimizer [56] to calculate the TD-error and perform backpropagations to update the network weights. Additionally, the Rainbow agents use distributional reinforcement learning to predict the value distributions which are approximated as a discrete distribution over 51 uniform atoms, along with prioritise replay memory sampling [29]. Even though Rainbow uses n-step bootstrapping, in these experiments a value of n=1 was found to be optimal, additionally noisy nets have been disabled in favour for a traditional decaying epsilon-greedy approach.
@@ 5.2 Self-play Performance
The learning curves for Rainbow with and without an augmented action-convention space for 2–5 player Hanabi are shown in Fig. 4. The two-player scenario seen in Fig. 4(a) shows the smallest improvement when applying conventions, however the two-player Hanabi problem can be considered a special case. This is due to the fact that the shared information between the two players are far more limited when compared to higher player counts. For example, in two-player Hanabi, half of the hands are hidden to a single player whereas in the three-player scenario only a third of the hands are hidden, and the players always have a common hand visible. This allows for more advanced reasoning over another player’s actions, evident by the fact that two of the conventions, specifically the Finesse and the Prompt as discussed in Appendix A.2, cannot be applied in two-player Hanabi.
(a) 2-Player learning curves
(b) 3-Player learning curves
(c) 4-Player learning curves
(d) 5-Player learning curves
Figure 4: Learning curves with exponential moving averages (weight=0.9995) for Rainbow, obtained from Bard et al. [37], as baseline compared to Rainbow with an augmented action-convention space for Hanabi two-to five-players. The best agent is highlighted in black for each agent scenario within each player count.
Despite the fact that the complexity of the problem increases exponentially as more agents are added to the environment, the performance uplift of conventions become more apparent. In the three-player setting seen in Fig. 4(b), the agents are able to achieve a significantly faster training time as well as an improved convergent performance with a higher average score. The performance improvement becomes even more apparent when looking at the five-player setting, seen in Fig. 4(d), which is generally considered the most difficult problem to solve, and where the agents train roughly five times faster than the baseline Rainbow agents. In Fig. 5 the distribution of scores during evaluation of each agent over the course of 1000 episodes is shown. In each scenario, conventions allow for a significant performance improvement, achieving a consistent grouping near a score of 21/25 and less overall variance when compared to baseline Rainbow. Rainbow with an augmented action-convention space also displays a robustness to increased player counts, with a remarkably consistent behaviour across all scenarios.
(a) 2-Player
(b) 3-Player
(c) 4-Player
(d) 5-Player
@@
Table 1 shows the average score for the best agent from each player count’s various runs compared to the results from Bard et al. [37]. Rainbow with an augmented action-convention space demonstrates an improved score over baseline Rainbow for each scenario, with a higher average in the 3–5 player scenario and overall less deviation. Additionally, the three-and five-player results are competitive compared to other algorithms, such as ACHA, while also requiring substantially less training steps. In the case of five-player Hanabi, Rainbow with conventions outperform ACHA where ACHA required 20 billion training steps and Rainbow with conventions only required 30 million. A statistical significance test, conducted in Appendix C, shows that majority of the results are statistically different. However, the Rainbow with conventions agent’s scores are statistically similar to that of baseline Rainbow two-players and ACHA three-players, however conventions allow the Rainbow algorithm to reach these scores significantly faster (5x compared to baseline Rainbow and 1000x compared to ACHA). Furthermore, even though these two instances are statistically similar, the agents learn completely different policies, as seen by the cross-play performance evaluation in the next section.
@@ 5.3 Cross-play Performance
Table 2 shows the two-player performance for various agents in cross-play Hanabi taken from existing literature, compared to the results of our cross-play Rainbow agents with an augmented action-convention space. The results for the Rainbow with conventions agents are statistically different from the other results reported in Table 2, as seen by Table 7 in Appendix C. Conventions are able to significantly improve on the performance of cross-play agents, allowing the baseline Rainbow agents to go from abysmal performance to an acceptable score. Furthermore, it is able to outperform SAD with only other-play applied for the two-player scenario, with other-play and auxiliary task applied to SAD still holding the best score. However, as mentioned previously, auxiliary tasks only benefit the two-player scenario in Hanabi [13]. When moving from self-play to cross-play, Rainbow with an augmented action-convention space only suffered a slight performance decrease and maintained remarkable consistency across various agent parings.
Table 2: Performance evaluation in two-player Hanabi for the best performing agents in a cross-play scenario over the course of 1000 episodes. The Rainbow with an augmented action-convention space agents were chosen from the top performing two agents in 10 different runs. Each score is averaged out of 25 with the standard error of the mean shown in brackets.
@@
Table 3 shows the results for cross-play Rainbow with an augmented action-convention space for all player counts of Hanabi. As seen in the self-play results for Rainbow with conventions and confirmed in Table 3, the two-player scenario is a special case, with the agents showing the biggest performance degradation when moving from self-play to cross-play, and overall worst score compared to the other cross-play results. While most research efforts on cross-play Hanabi focus on the two-player scenario, our results demonstrate that more powerful conventions are applicable in higher player counts, thus, it is important to consider all the player scenarios when evaluating cross-play performance.
@@
Most remarkable is the five-player results, where the cross-play Rainbow agents with an augmented action-convention space are able to achieve a relatively high score and still cooperate effectively. The reason conventions are able to offer such a large performance gain for cross-play is due to the fact that the agents learn from a communal list of conventions that are taken from existing domain knowledge. Thus, if an agent is paired with a never-before-seen cooperator that has learned to play with the same list of conventions, they will still be able to cooperate effectively, since there is no uncertainty over another agent’s intent and reasoning. This is not unexpected, since humans who learn from the same “list” of conventions (whether it is physical or passed down), and that have never before interacted, can cooperate and communicate effectively with relatively low uncertainty.
@@ 6 Conclusion
Our main results show that conventions are able to significantly improve on the training time for Rainbow agents in Hanabi, and in the case of three-to five-players, also improve on the converged policy. Other Hanabi algorithms, such as SAD or MAPPO, require billions of training steps to reach convergent policies, whereas conventions reduced the number of training steps for Rainbow to below 30 million consistently. Additionally, conventions showed significant performance increase for cross-play agents, and allowed the Rainbow agents to go from not cooperating at all to being able to achieve decent performance across all player scenarios. For the most difficult scenario of five players, the cross-play Rainbow conventions agents trained on 50 million steps are able to outperform the ACHA self-play agents trained on 20 billion steps. Our work, as presented in this paper, mainly focused on single-and two-step conventions in the Hanabi problem, with the goal of extending these concepts to other multi-agent scenarios (including simultaneous action spaces) containing multistep conventions, as well as the exploration of more advance conventions (and the application of those conventions) on more advanced algorithms within the Hanabi setting.